OpenStudy (anonymous):

Lyle purchased a motorcycle for $13,430. It depreciates about 4.1% each year. What is the value of the motorcycle after seven years? $13,112.07 $9,575.59 $10,018.58 $17,792.26

OpenStudy (phi):

it goes down by 4.1% or 0.041 per year that means it has a value of (1-0.041)= 0.959 of the previous year try 0.959^n * 13430 where n is the number of years

OpenStudy (anonymous):

Thanks.. So I replaced n with 7 and got 10,018.58... Is that accurate?

OpenStudy (phi):

yes. Does the idea make sense? after one year it is worth 0.959*13430 after the next year, it goes down by another factor of 0.959 or 0.959*0.959*13430 = 0.959^2 * 13430 we keep doing that for 7 years.

OpenStudy (anonymous):

So the only thing that would be altered would be n for years.. Alright. Gotcha. Thanks.